# Simulator Ensembles for Trustworthy Autonomous Driving Testing

### v1

@@ 5.7.2 Efficiency (RQ2)
The efficiency results are shown in Table 3. The results show the ratio of the search budget and its standard deviation when the first valid failure is detected. We observe that all MultiSim configurations exhibit higher percentages for the first valid failure compared to SingleSim, as expected. Compared to DSS, we see that the configuration BD achieves a better result than DSS-BD. However, across all configurations, the standard deviations for MultiSim are higher than those of SingleSim or DSS. The combination UD is surpassed by SingleSim and DSS approaches, beside DSS-BD which yields a higher search time. For BU we can observe a similar behavior.
The statistical test results are shown in Table 4. While D significantly outperforms DB, DB significantly performs better then DSS-DB. Regarding BU, we can only observe that it is outperformed by DSS-BU and D, while there is no significant difference with the other approaches. The configuration UD yields also significant lower search times then U and D, with large effect sizes. However, we do not observe a statistical significant difference w.r.t. DSS-UD, which confirms our observations from Table 4.
Table 3: Efficiency results, i.e., time to first valid failure (in percentage w.r.t. the search budget).
Method Ratio Budget Std Dev
BD 39.7 22.9
UD 35.3 24.0
BU 55.6 29.4
U 22.8 16.1
B 25.5 11.3
D 22.4 14.1
DSS-BD 56.6 16.3
DSS-UD 39.9 13.6
DSS-BU 33.5 13.2
Table 4: Efficiency. Statistical tests (Wilcoxon and Vargha-Delaney) for identifying the first valid failure (repeated comparisons are excluded (N/A)). Effect measure values are annotated with a star, if the first approach of the comparison yields lower values then the compared approach.
@@
0.03 (L) 0.02 (L) 0.01 (L) N/A N/A 0.01 (L*) 0.92 (-) 0.49 (-)
Figure 6: Visualization of the tradeoff between efficiency (first valid failure) and effectiveness (valid_rate) for MultiSim, DSS and SingleSim for different configurations. The Pareto front is visualized in dashed red. Validity rate values are negated, to visualize the Pareto front in the bottom left part of the Figure.
@@
Figure 6 shows the study results considering both effectiveness and efficiency (validity rate vs. search budget required for first valid failure) for MultiSim, SingleSim and DSS. Particuarly, the methods whose values are non-dominated either in effectiveness nor efficiency are BD, B and D (Pareto-optimal solutions). SingleSim, when using BeamNG and Donkey, is more efficient in finding the first valid failure, while based on the results of the MultiSim combination BD, MultiSim is significantly more effective.
RQ2 (efficiency): In two out of three simulator configurations (BD and BU), MultiSim demonstrates efficiency comparable to SingleSim. Compared to multi-simulator based testing, in the BD configuration, MultiSim achieves significantly better results than DSS. However, when considering BU, DSS-BU outperforms MultiSim.
5.7.3 Prediction RQ3
We selected the best performing MultiSim configuration, i.e., BD, to be extended to support prediction of disagreements. In particular, we selected all test cases from 5 completed runs with BD and 5 additional search runs with a modified version of MultiSim that focuses on disagreements (Section 5.3). In total, we collected a dataset of 580 agreements and 572 disagreements, which we used to train 5 classification models in predicting disagreements.
The results of the 5-fold cross-validation, using an 80:20 training-to-test split for the trained classifiers, are presented in Table 5. DT represents Decision Tree, RF represents Random Forests, SVM represents Support Vector Machine, LREG stands for Logistic Regression, and GBOOST stands for Gradient Boosing. From all trained classifiers, we select the best performing classification model according to the metrics of Accuracy, F-1 Score and AUC. We identify the RF model as the best classifier and integrate it into MultiSim to predict disagreements as described in Section 5.3. As uncertainty threshold we select 70%.
Effectiveness. The boxplots for the valid_rate and n_valid metrics averaged over 10 runs are shown in Figure 7. The corresponding statistical test results are as follows: regarding the n_valid metric, we identify, for the comparison BD vs. BD-P (i.e., the MultiSim version of the BD configuration but with the disagreements predictor), a p-value of 0.046 with a large effect size. While the difference is significant, we can see from Figure 7 that the median valid_rate is still above 90%. As for n_valid, we do not observe a statistically significant difference as the comparison yields a p-value of 0.76 with a medium effect size. However, from Figure 7 we see that the median value is by 5 failures higher than that of BD-P.
Figure 7: Effectiveness (a,b) and Efficiency results (c) of MultiSim with disagreements prediction averaged over 10 runs. BD-P indicates MultiSim with disagreements prediction while BD indicates the original MultiSim approach.
@@ 5.7.3 Prediction RQ3
Table 5: Classifier results after 5-fold cross-validation in percentages for predicting disagreements in the simulator configuration BD. Average values are reported with standard deviations in brackets. The best results are marked in bold.
Model Accuracy Precision F1 Score AUC
DT 78 (2) 78 (4) 79 (1) 78 (2)
RF 83 (1) 82 (3) 83 (1) 89 (2)
SVM 62 (2) 58 (2) 69 (2) 71 (4)
LREG 58 (4) 58 (5) 57 (4) 62 (4)
GBOOST 80 (3) 78 (3) 80 (3) 87 (3)
Efficiency. The boxplot results comparing the efficiency of BD-P and BD are shown in 7(c). BD-P requires on average less time, i.e., 33% of the search budget compared to BD that finds the first valid failures after 39% of the available search budget. Additionally, we can observe that the standard deviation of BD-P is smaller (14.2%) compared to that of BD (22.9%). However, the statistical test results yield a p-value of 0.49, which shows no significant differences between the two approaches.
RQ3 (prediction): Among all trained classification models for predicting simulator-related disagreements in simulation outcomes, Random Forest performs best, achieving an AUC score of 0.89 and an F1 score of 0.83. Integrating a classification model into MultiSim to bypass simulations leads to a higher median, average and lower standard deviation of simulator-agnostic failures compared to conducting the search without the disagreements predictor.
@@ 5.8.1 Disagreements
We manually analyzed the disagreements found by MultiSim across different configurations. We made following observations:
(Observation 1) In the BeamNG simulator, occasionally, the vehicle switches to the left lane whenever the underlying segment was short, i.e., $<$ 15m, and the following segment was a left turn (Figure 8). The car continues driving on the left lane towards the next segment, incurring in a failure. This behaviour could be explained by the fact that the vehicle observes the upcoming segment, and considers this shortcut as the immediate straight continuation of its path. However, in Donkey, the vehicle stays in the right lane, even when driving on the same short segment.
(Observation 2) In the BeamNG simulator, the vehicle switches to the left lane in the last but one segment after meeting a sharp curve. This behaviour might be related to the fact that the road terminates after the last segment. However, we did not observe a similar behaviour for the vehicle in the Donkey or Udacity Simulator, which might be attributed to the visualization/rendering differences in the simulators.
@@
Figure 9: Feature map of tests found in one run with MultiSim employing disagreement classifier (BD-P). Cells containing roads for which disagreements are predicted are marked in orange with a black box.
@@ 5.8.2 Feature Maps of Prediction-based Search
When evaluating the results of MultiSim augmented with the prediction classifier, we inspected the feature maps before failure validation. We could observe that disagreements are predicted in regions which are between failing and non-failing cells of the map, as shown in Figure 9. Cells that contain tests that are predicted as disagreements by the classifier are marked as orange. This observation highlights the fact that our classifier is likely to perform accurately in predicting disagreements, as it is predicting disagreement on test cases which lie on the boundary between scenarios/roads that pass and fail.
@@ 6 Discussion
Approach. While our study utilizes two simulators in the ensemble, MultiSim can be extended to incorporate a larger number of simulators. To reach consensus on the criticality, we suggest to use an odd number of simulators. Further, the feasibility of the approach with multiple simulators is basically constrained by the underlying computational resources. To mitigate this, test executions in MultiSim with multiple simulators might be parallelized, independent of the underlying test algorithm used.
Lessons Learned. Our case study shows that employing multiple simulators in the search does improve the effectiveness of finding simulator-agnostic failures compared to single-simulator testing. However, in the literature, test case generation approaches in general report the evaluation results which are based on single-simulator executions. We suggest that failures found during the test generation or at least the final identified failures, should be evaluated in a second simulator to assess whether the failures are likely to be simulator-agnostic.
Root Cause Analysis. Our paper proposes a testing technique to identify failing test cases where simulators agree on. It is out of the scope of this study to investigate the underlying root causes of disagreements between simulator executions, as it would likely require to inspect, debug, and profile the simulator’s code (in Section 5.8 we qualitatively analyzed two disagreement scenarios). Additionally, as outlined in RQ3, the MultiSim approach can be configured to search for tests on which simulators disagree to support the identification of the root causes behind failures. As shown by Jodat et al. JodatSpurious23 and Ben Abdessalem et al. Abdessalem-ICSE18 , decision trees or decision rules can be then applied on underlying roads and simulation traces to derive conditions on input variables and analyze the root causes behind inconclusive and simulator-specific failures.
@@
Validation. In our study, the validation was dependent on three hyperparameters: the number of re-executions, the failure rate threshold, and the number of failures selected per cell. We set the failure rate threshold to 100%. However, in practice, the threshold for the validation can be configured based on the number of simulators used. We expect that the more diverse simulation environments are employed, the more difficult would be to achieve failure rates of 100%. Regarding the number of re-executions we suggest to perform preliminary experiments with a different number of re-executions and compare the results, as conducted in our study. To select the number of selected failing tests per cell, we carried out preliminary experiments with a higher number of failures, i.e., 5. However, we did not observe significantly different evaluation results. Moreover, we did not select more tests per cell because the feature map already discretizes the search domain, making the diversity of roads within a single cell being likely low.
@@ 7 Threats to Validity
External validity is related to the generalizability of our results. We have evaluated MultiSim on a well known case study, a DNN trained for lane keeping, employing different simulator pairs from a selected set of different simulators (BeamNG, Donkey and Udacity). It would be interesting to analyze whether the results apply for other case studies, e.g., including automated emergency braking system in scenarios where other actors are deployed. However, aligning the scenarios, visual appearance, and simulation conditions across simulator with other actors poses a challenge compared to the lane-keeping ADAS case study. Another threat is related to the usage of one ADAS in our evaluation, given its high computational cost (more than 22 days for test generation and validation). Results might not transfer to other ADAS models. However, DAVE-2 is a popular and reliable model across virtual and real-world testing 2023-Stocco-EMSE ; 2023-Stocco-TSE , and we mitigated the threats related to an insufficiently trained models with our preliminary evaluation.
@@
Internal validity risks refer to confounding factors. To mitigate confounding factors, we have taken several mitigation strategies. First, to avoid a biased evaluation because of unaligned deployment conditions, i.e., major deviations in communication delays between the SUT and the simulator, we have performed a preliminary assessment/validation of the XTE variation for manually defined driving scenarios across the simulators. Second, we made sure that scenario/road simulation is aligned across simulators building our study upon existing research matteoMaxibon2024 . Third, we have defined a condition when generated roads are considered as invalid, and have regenerated after mutation roads when they are invalid.
To evaluate the BU setup, we have executed each simulator on different hardware configurations, because of simulator-system incompatibilities (Udacity did not run on Windows, but BeamNG only runs on Windows). This fact might have influenced BU results. However, the BU configuration did not yield a superior performance compared to approaches highlighted in our results.
Further, to make sure that we identify simulator-agnostic failures we validated identified failures after the test generation process of MultiSim, DSS and SingleSim on one or two simulators which has been left-out during the search. In addition we re-ran failing tests multiple times to filter out failures identified because of a flaky simulation environment.
Construct validity threats relate to the inappropriate use of metrics. To evaluate the performance of MultiSim and compared approaches we have used two metrics: the number of simulator-agnostic failures, and the validity rate, i.e., the ratio between simulator-agnostic failures and validated failures. The first metric is a default metric used to assess the performance of test case generation approaches Kluck19Nsga2ADAS . The choice of the second metric is motivated by the fact that, when using a test generator, the tester is interested in knowing how many of the generated failures are simulator-agnostic. In addition, we analyzed the convergence of the HyperVolume metric applied to MultiSim and SingleSim, which indicates that running longer SingleSim search will not improve effectiveness results. We did not evaluate the convergence of DSS as it is not a Pareto-based optimization approach.
To foster replicability, the implementation of MultiSim, as well as results of the evaluation for MultiSim, SingleSim and DSS are available online. The search approach and the analysis of the results have been implemented using the modular and open-source search-based testing framework OpenSBT sorokin2023opensbt , which allows to integrate other case studies to be tested with MultiSim, SingleSim and DSS.

### v2

@@ 5.7.2 Efficiency (RQ2)
Table 3 shows the efficiency results, specifically the ratio of the search budget and its standard deviation when the first valid failure is detected.
For DAVE-2 we observe that all MultiSim configurations exhibit higher percentages for the first valid failure compared to SingleSim, as expected. Compared to DSS, we see that the configuration BD achieves a better result than DSS-BD. However, across all configurations, the standard deviations for MultiSim are higher than those of SingleSim or DSS. The combination UD is surpassed by SingleSim and DSS approaches, beside DSS-BD which yields a higher search time. For BU we can observe a similar behavior.
The statistical test results are shown in Table 4. While for DAVE-2 D significantly outperforms BD, BD significantly performs better then DSS-BD. Regarding BU, we can only observe that it is outperformed by DSS-BU and D, while there is no significant difference with the other approaches. The SingleSim configurations U and D yield significantly lower search times then UD, with large effect sizes. However, we do not observe a statistical significant difference of UD w.r.t. DSS-UD, which confirms our observations from Table 4.
For ViT, we can see that, similarly as for DAVE-2, MultiSim combinations achieve on average higher values than SingleSim-based search and DSS.
For TCP, MultiSim achieves lower scores with similar variation compared to SingleSim. However the difference is not statistically significant except for the comparison between BU and B for TCP, where MultiSim achieves significantly better results.
Table 3: Efficiency results, i.e., time to first valid failure for DAVE-2, ViT, and TCP. Mean and standard deviation (std) are in per cent.
Method DAVE-2 ViT TCP
Mean Std Mean Std Mean Std
BD 39.7 22.9 34.2 9.4 40.6 31.2
UD 35.3 24.0 56.5 22.2 31.0 33.3
BU 55.6 29.4 60.1 29.4 30.6 31.2
U 22.8 16.1 44.5 37.3 59.0 31.1
B 25.5 11.3 51.4 31.1 79.5 26.9
D 22.4 14.1 37.0 20.9 74.4 34.4
DSS-BD 56.6 16.3 37.1 2.2 61.1 12.5
DSS-UD 39.9 13.6 43.4 6.5 45.8 63.5
DSS-BU 33.5 13.2 34.6 5.2 51.6 10.2
(a) DAVE-2
(b) ViT
(c) TCP
Figure 6: Visualization of the tradeoff between efficiency (first valid failure) and effectiveness (valid_rate) for MultiSim, DSS and SingleSim for different configurations across three case studies. The Pareto front/point is visualized in red.
Table 4: (DAVE-2) Efficiency. Statistical tests (Wilcoxon and Vargha-Delaney) for identifying the first valid failure (repeated comparisons are excluded (N/A)) for the ADAS DAVE-2, ViT and TCP. Effect size magnitudes are annotated with a star, if the first approach (rows) of the comparison yields lower values than the second approach (columns).
@@
0.03 (L*) 0.02 (L*) 0.01 (L*) N/A N/A 0.01 (L*) 0.92 (-) 0.49 (-)
ViT
BD
1.00 (L) 0.31 (S) 0.84 (S) 0.16 (L*) 0.16 (L) 0.06 (L) 0.84 (-) 0.09 (L)
BU
0.56 (M) 0.56 (S) 0.44 (L) N/A 0.69 (-) 0.09 (L) 0.16 (L) 0.31 (M)
UD
0.56 (M) 0.84 (S) 0.31 (L) N/A N/A 0.16 (L) 0.16 (L) 0.44 (S)
TCP
BD
0.22 (M*) 0.16 (L*) 0.09 (L*) 0.31 (L) 0.44 (L) 0.31 (L*) 0.56 (L*) 0.56 (L*)
BU
0.16 (L*) 0.04 (L*) 0.16 (L*) N/A N/A 0.09 (L*) 0.31 (L*) 0.44 (L*)
UD
0.22 (L*) 0.16 (L*) 0.16 (L*) N/A 1.00 (-) 0.06 (L*) 0.44 (L*) 0.44 (L*)
@@
Figure 6 shows the study results considering both effectiveness and efficiency (validity rate vs. search budget required for first valid failure) for DAVE-2, ViT and TCP for MultiSim, SingleSim and DSS. Particuarly, for DAVE-2 the methods whose values are non-dominated either in effectiveness nor efficiency are BD, B and D (Pareto-optimal solutions). SingleSim, when using BeamNG and Donkey, is more efficient in finding the first valid failure, while based on the results of the MultiSim combination BD, MultiSim is significantly more effective. For ViT/TCP, we see that the MultiSim combination BD/BU is dominating all remaining configurations yielding both a high validity rate with a low search effort for the first valid failure.
$RQ_{2}$ (efficiency). Across all configurations, MultiSim achieves efficiency comparable to SingleSim, requiring a similar search budget to find the first valid failure. Among multi-simulator settings, BD and UD are the most efficient for DAVE-2 and ViT, while BU performs best for TCP. Compared to DSS, in general MultiSim shows better efficiency, whereas DSS-BU remains in two out of three comparisons slightly more time-efficient than BU. Overall, MultiSim maintains competitive efficiency while yielding substantially higher failure validity.
5.7.3 Prediction (RQ3)
To reduce unnecessary simulator evaluations, we extend MultiSim with a surrogate disagreement predictor. This surrogate is a classifier, trained using labeled outcomes of previous runs, where each instance is annotated as an agreement or disagreement. The surrogate outputs the probability of a disagreement event between simulator pairs. Specifically, we gathered all test cases from five completed runs using BD, along with five additional runs using a modified version of MultiSim focused on identifying disagreements (Section 5.3). In total, we collected a dataset comprising 580 agreements and 572 disagreements for DAVE-2, 581 agreements and 581 disagreements for ViT, and 383 agreements and 383 disagreements for TCP. These datasets were then used to train five classification models aimed at predicting disagreement occurrences. We train and validate multiple surrogate classifiers (Decision Tree, Random Forest, SVM, Logistic Regression, and Gradient Boosting) using 5-fold cross-validation with an 80:20 split. As reported in Table 5, the Random Forest model achieved the best performance for every system under test, with F-1 scores close to or above 80%. We therefore integrate this model into MultiSim as the surrogate predictor.
During the test generation phase, each newly generated candidate scenario is first evaluated by the surrogate model. If the predicted probability of simulator disagreement exceeds a threshold ($\tau=0.7$), the candidate is discarded before running the full simulation. Otherwise, it is executed across both simulators. This filtering mechanism allows MultiSim to save simulation budget by focusing exploration on high-value, disagreement-prone regions of the search space.
Across all three case studies, integrating the surrogate disagreement predictor into MultiSim (BD-P) improved the overall test generation effectiveness. The surrogate-guided runs produced a higher number of valid and fault-revealing test cases compared to the baseline configuration (BD), increasing the mean n_valid and valid_rate by approximately 38% on average. This confirms that the surrogate model not only reduces redundant simulator queries but also helps focusing the search on higher-yield regions of the input space, maintaining an average validity rate above 90% across runs.
The use of the surrogate model also leads to efficiency improvements in two out of the three case studies. For DAVE-2 it yields an improvement of 14%, and 2% for TCP. These results indicate that surrogate-based filtering can substantially accelerate simulation-based testing while maintaining the ability to identify a broad range of failure cases.
@@ 5.7.3 Prediction (RQ3)
Table 5: Classifier results after 5-fold cross-validation (in %) for predicting disagreements in three simulator configurations. The best results are marked in bold.
DAVE-2 ViT TCP
F-1 AUC-ROC F-1 AUC-ROC F-1 AUC-ROC
DT 79 78 58 65 60 65
RF 83 89 74 81 78 86
SVM 69 71 65 77 58 80
LREG 57 62 58 65 60 65
GBOOST 80 87 73 80 75 82
Table 6: Mean and standard deviation of effectiveness and efficiency metrics for BD-P and BD approaches across three use cases, averaged over 10 runs for DAVE-2, 6 runs for ViT, and 6 runs for TCP.
Metric DAVE-2 ViT TCP
BD-P BD BD-P BD BD-P BD
n_valid 25.3 $\pm$ 7.3 23.0 $\pm$ 13.1 15.8 $\pm$ 2.0 15.5 $\pm$ 8.3 8.3 $\pm$ 3.6 4.1 $\pm$ 4.8
valid_rate (%) 92.9 $\pm$ 7.9 98.5 $\pm$ 3.0 97.9 $\pm$ 3.0 98.1 $\pm$ 4.1 44.3 $\pm$ 8.3 25.9 $\pm$ 18.7
first_fail (%) 34.0 $\pm$ 14.1 39.7 $\pm$ 22.9 43. $\pm$ 16.7 34.2 $\pm$ 9.4 39.8 $\pm$ 25.9 40.6 $\pm$ 31.2
RQ3 (prediction). Among all trained classifiers, the Random Forest achieved the best disagreement-prediction performance (AUC: 0.81–0.89, F-1: 0.74–0.83). Integrating this surrogate into MultiSim enabled the framework to bypass low-value simulations, improving both effectiveness and efficiency by increasing the number of valid failures and reducing variability compared to baseline approaches.
@@ 5.8.1 Disagreements
We manually analyzed the disagreements found by MultiSim across different configurations. We made following observations.
In the BeamNG simulator for DAVE-2 occasionally, the vehicle switches to the left lane whenever the underlying segment is short, i.e., << 15m, and the following segment is a left turn (Figure 7). The car continues driving on the left lane towards the next segment, incurring in a failure.
A similar behavior was observed for TCP in Donkey and Udacity for sharp and short left turns followed by a sharp right turn, where the vehicle did not continue driving to the left, but proceeded straight to align with the subsequent path after the right turn. This behaviour could be explained by the fact that the vehicle observes the upcoming segment, and considers this shortcut as the immediate straight continuation of its path. However, in the remaining simulators and system under test configurations the vehicle stays in the lane, even when driving on short segments.
Specifically to TCP, we observed that the model exhibits partially in all simulations environments a slightly higher XTE variation compared to DAVE-2 and ViT even when the single-simulator based failure count is less then for DAVE-2 or ViT. At the same time we could see that the model steers in general more extreme in curves compared to DAVE-2 or ViT which could explain its robust performance and the lower failure rate compared to the other system under tests.
Across all system under tests, we observed also that in the BeamNG simulator, the vehicle switches often to the left lane in the last but one segment after meeting a sharp curve. This behaviour might be related to the fact that the road terminates after the last segment. However, we did not observe a similar behaviour for the vehicle in the Donkey or Udacity simulators, which might be attributed to the visualization/rendering differences in the simulators.
@@
Figure 8: Feature map of tests found in one run with MultiSim employing disagreement classifier (BD-P) for DAVE-2. Cells containing roads for which disagreements are predicted are marked in orange with a black box.
@@ 5.8.2 Feature Maps of Prediction-based Search
When evaluating the results of MultiSim augmented with the prediction classifier, we inspected the feature maps before failure validation. We could observe that disagreements are predicted in regions which are between failing and non-failing cells of the map. An example for DAVE-2 is shown in Figure 8. Cells that contain tests that are predicted as disagreements by the classifier are marked as orange. This observation highlights the fact that our classifier is likely to perform accurately in predicting disagreements, as it is predicting disagreement on test cases which lie on the boundary between scenarios/roads that pass and fail. Similar situations were observed in the other simulator ensembles.
@@ 6 Discussion
6.1 Approach
While our study utilizes two simulators in the ensemble, MultiSim can be extended to incorporate a larger number of simulators. To reach consensus on the criticality, we suggest to use an odd number of simulators. Further, the feasibility of the approach with multiple simulators is basically constrained by the underlying computational resources. To mitigate this, test executions in MultiSim with multiple simulators might be parallelized, independent of the underlying test algorithm used.
6.2 Lessons Learned
Our case study shows that employing multiple simulators in the search does improve the effectiveness of finding simulator-agnostic failures compared to single-simulator testing. However, in the literature, test case generation approaches in general report the evaluation results which are based on single-simulator executions. We suggest that failures found during the test generation or at least the final identified failures, should be evaluated in a second simulator to assess whether the failures are likely to be simulator-agnostic.
6.3 Root Cause Analysis
Our paper proposes a testing technique to identify failing test cases where simulators agree on. It is out of the scope of this study to investigate the underlying root causes of disagreements between simulator executions, as it would likely require to inspect, debug, and profile the simulator’s code (in Section 5.8 we qualitatively analyzed two disagreement scenarios). Additionally, as outlined in RQ3, the MultiSim approach can be configured to search for tests on which simulators disagree to support the identification of the root causes behind failures. As shown by Jodat et al. JodatSpurious23 and Ben Abdessalem et al. Abdessalem-ICSE18, decision trees or decision rules can be then applied on underlying roads and simulation traces to derive conditions on input variables and analyze the root causes behind inconclusive and simulator-specific failures.
@@ 6.3 Root Cause Analysis
We observed that BD performs best for DAVE-2 and ViT, whereas BU excels for TCP. This can be explained by the differing sensitivities of the ADAS architectures to physical versus perceptual variability. BD combines simulators with heterogeneous vehicle dynamics (soft- vs. rigid-body physics), which better exposes control-related faults, while BU combines simulators with distinct rendering and camera models, thus better revealing perception-related inconsistencies. Consequently, it appears that the optimal simulator ensemble depends on the dominant failure modes of the SUT.
6.4 Validation
In our study, the validation was dependent on three hyperparameters: the number of re-executions, the failure rate threshold, and the number of failures selected per cell. We set the failure rate threshold to 100%. However, in practice, the threshold for the validation can be configured based on the number of simulators used. We expect that the more diverse simulation environments are employed, the more difficult would be to achieve failure rates of 100%. Regarding the number of re-executions we suggest to perform preliminary experiments with a different number of re-executions and compare the results, as conducted in our study. To select the number of selected failing tests per cell, we carried out preliminary experiments with a higher number of failures, i.e., 5. However, we did not observe significantly different evaluation results. Moreover, we did not select more tests per cell because the feature map already discretizes the search domain, making the diversity of roads within a single cell being likely low.
6.5 Generalization of Our Approach
While our empirical evaluation focuses on lane-keeping assistance systems, the MultiSim framework is not limited to this specific ADAS function. The proposed architecture and workflow, based on simulator coupling, multi-objective search, and cross-simulator agreement analysis, can be generalized to other ADAS and ADS functionalities.
First, the representation of tests is model-based and modular: road geometry is encoded through control points and spline interpolation, but this abstraction can be extended to other types of driving scenarios. For instance, intersections, pedestrian crossings, or obstacle avoidance can be represented using higher-level scene graphs or semantic maps instead of simple lane curves. Similarly, perception-heavy ADAS such as automatic emergency braking (AEB), adaptive cruise control (ACC), or lane-change assist could be evaluated by defining task-specific fitness functions (e.g., time-to-collision, minimum distance to obstacles, or stability of steering trajectories).
Second, the optimization procedure is independent of the specific fitness formulation. Our search and repopulation strategies can operate over any domain where simulation feedback is measurable, making it possible to reuse the same search backbone for tasks with different control policies or sensing modalities (e.g., radar-, LiDAR-, or camera-based). This in particular is enabled by the modular architecture of the extended framework OpenSBT as shown in various case studies sorokin2023opensbt.
Finally, the MultiSim framework could be extended to support system-level ADS validation by incorporating full autonomy stacks like Autoware autoware or Apollo apollo. In this context, cross-simulator disagreement analysis could help identify environment- or simulator-induced discrepancies in perception, planning, and control, offering a systematic means of analyzing simulation fidelity and testing robustness of the end-to-end driving system. However, since no existing autonomous driving stack is truly simulator-agnostic, extending MultiSim to full ADS validation remains an open engineering challenge that requires standardized interfaces and consistent scene representations across simulators.
@@ 7 Threats to Validity
External validity concerns the extent to which our results can be generalized beyond the studied settings. We evaluated MultiSim on a well-established case study involving three lane-keeping models of increasing complexity, each tested across different simulator pairs drawn from a diverse set of environments, BeamNG, DonkeyCar, and Udacity. It would be interesting to investigate whether the results generalize to other ADAS functions, such as automated emergency braking, or to higher levels of autonomy using full driving stacks like Autoware autoware or Apollo apollo. This would require testing in more complex urban simulation environments, such as AWSIM awsim or CARLA carla. The main difficulty lies in the technical integration of modern autonomous driving stacks with multiple simulators, as these frameworks are typically coupled with a single specific simulator. Achieving cross-simulator consistency in terms of scenarios, visual appearance, simulation conditions, and interactions with other actors poses significantly greater challenges than in our current study.
@@
Internal validity risks refer to confounding factors that could affect the interpretation of results. To mitigate these, we adopted several strategies. First, to avoid biased evaluations due to unaligned deployment conditions (e.g., communication delays between the SUT and the simulator), we conducted a preliminary validation of XTE variation across manually defined driving scenarios. Second, we aligned the simulated roads across all simulators, building upon existing research matteoMaxibon2024. Third, we defined validity constraints for generated roads and regenerated invalid samples after mutation. For the BU setup, simulators were executed on different hardware configurations due to compatibility constraints (BeamNG requires running on a Windows operating system). This may have influenced BU results, although BU did not outperform other configurations in our experiments. To ensure that detected failures were simulator-agnostic, we validated all failures identified by MultiSim, DSS, and SingleSim on one or two simulators left out during the search. We also re-ran failing tests multiple times to filter out failures caused by transient or flaky simulation conditions.
Construct validity threats relate to the adequacy of the chosen metrics. To evaluate performance, we used two primary metrics: the number of simulator-agnostic failures and the validity rate, i.e., the ratio between simulator-agnostic failures and validated failures. The first metric is a standard measure in test case generation studies Kluck19Nsga2ADAS, while the second captures the proportion of failures that remain consistent across simulators, reflecting robustness and transferability. Additionally, we analyzed the convergence of the Hypervolume (HV) indicator for MultiSim and SingleSim, showing that SingleSim converges earlier, while MultiSim continues improving slightly over time. We excluded DSS from HV analysis as it is not Pareto-based.
To foster replicability, we publicly release the implementation of MultiSim, along with all experimental data for MultiSim, SingleSim, and DSS. The search algorithms and result analyses are implemented using the modular and open-source framework OpenSBT sorokin2023opensbt, which supports easy integration of additional case studies for future research.